# HHardware specifications

### v1

@@ Abstract
footnotetext: $\dagger$ For all inquiries please contact: raja.yehia@icfo.net
1 Introduction
@@
With the end goal of constructing a global quantum internet [1], quantum networks are rapidly developing and are already entering a deployment phase. Several national and international initiatives [2, 3] are already in place to establish an architecture allowing for distant nodes to perform quantum cryptographic tasks, such as secret key exchange. In a world with finite resources where energy demands outgrow energy generation, it is therefore crucial to estimate how much energy these networks will consume prior to their deployment [4]. Such a study can reveal limiting factors for future implementations of networks, or even show the energetic advantages of certain quantum technologies over classical ones. Works that estimate the energetic cost of quantum devices are few, but there are indications [5, 6] that quantum computing may show an energetic advantage before a computational one.
@@
Comparing classical and quantum communication protocols in terms of energy is a difficult task, mainly because there are no classical equivalents to most quantum networking protocols with the same level of security. For example, a quantum key distribution protocol achieving information-theoretic security cannot adequately be compared to a classical communication protocol providing only computational security. It is possible, however, to compare different quantum protocols achieving the same functionality. The focus of this study will be put on this task in addition to benchmarking the effective energy consumption of near-term quantum communication infrastructures.
@@
This work presents the foundations of a framework to estimate the energy cost of quantum network protocols. A first estimation is given of the energy requirements of basic network functionalities, namely Quantum Key Distribution (QKD) and Conference Key Agreement (CKA), whose goals are to generate a secret private key among end users of a quantum network. The methods and hardware they use are generic to most protocols based on photonic implementations. In particular, the creation and sharing of entangled states among distant parties, believed to be the main goal of most quantum network architectures [7], are the building blocks of many other network protocols [8, 9, 10, 11, 12, 13].
@@
To obtain concrete figures of merit, we take a hardware-dependent approach to compare different implementations of some common protocols. Namely, different QKD implementations are compared, and the implementation of networks of $N$ nodes are analyzed, since their scaling in resources with the number of nodes is non-trivial. Using the energetic cost as a benchmark, instead of the rate or the fidelity, gives a unique perspective. For example, our simulations suggest that there exists regimes of parameters for QKD protocols where using less efficient but more energy effective detectors results in huge energy savings at the cost of increased execution time. Another example of results from this work are the discovery of distance regimes for which the usage of different wavelengths results in energy savings, and the identification of optimal protocols to achieve multipartite tasks as a function of the number of parties.
@@
This work is organised as follows: first, the model for estimating the energy consumption of photonic setups is introduced in Section 2. The estimation of the energy costs of bipartite QKD protocols is presented in Section 3 and that of multipartite network protocols, in Section 4. To improve readability, the table of hardware components that was used as reference is deferred to Appendix A and the description of the modelling of their energy consumption, to Appendix B. Appendix C contains the description of the QKD protocols that were simulated, and additional results on time-bin based setups are presented in Appendix D, Appendix E displays experimentally measured energy values, and Appendix F touches on the distribution of the power consumption between the users.
2 Energy model
@@
The energy cost of any quantum network protocol can be divided into multiple contributions, which can be studied independently to estimate the total energy cost of the protocol. The first contribution is the source, a broad term encompassing all the required components to generate the state of light onto which the information is encoded. The second contribution to the energy cost regroups all manipulations required in the protocol which include, for example, changes in polarization applied through motorized waveplates, phase shifters, and so on. The third is the detection where the optical signal is measured. Finally, the last contribution comes from classical communications and computations that are required by the protocol.
@@
We define the total energy cost for a protocol $\pi$ as the function $E_{\pi}(t)$ which depends on time, and is composed of:
$E_{\pi}(t)=$ $\sum^{n_{S}}_{i}S^{i}_{\pi}(t)+\sum^{n_{M}}_{k}M^{k}_{\pi}(t)+\sum^{n_{D}}_{j}D^{j}_{\pi}(t)+C_{\pi}(t)+E_{\pi}(0).$ (1)
@@
On top of optical components, quantum network protocols typically include some classical computing elements that increase the energy cost, and those are denoted as $C_{\pi}(t)$. Two major families of contributions are included in this term. The first one consists of the classical controls of quantum components during the time of the protocol. Here, the costs of different components that are commonly involved in communication protocols are considered, such as time taggers to record timing events. All other classical costs required during a protocol are modelled by considering that an active computer is present at each node involved in the protocol. This covers classical communications and memory costs, or classical sub-protocols such as coin flipping.
The second family of contributions is the energy required for post-processing, which refers to the classical algorithms applied to the outcome of quantum processes. This cost is harder to estimate as it depends on various parameters such as the desired level of security, the classical computing architecture and the choice of post-processing functions that are usually not standardized. In this work, this cost is mostly overlooked. However, this post-processing cost can prove to be non-trivial for some of the protocols studied in this work, especially when considering digital signal processing. This particular cost is discussed in Section 3.3.
Each of these families of elements are expanded in Appendix B. In Table 3, some values are shown of power consumption for different types of devices that are used in the following sections.
3 Energy cost of bipartite protocols: Quantum Key Distribution
In this section, we investigate the energetic cost of different widespread QKD protocols. Specifically, the model presented in this work is tested with the well-known BB84 protocol proposed by Bennett and Brassard [14, 15, 16]. We also explore its entanglement-based version, the E91 protocol [17, 18], along with a measurement-device-independent QKD (MDI-QKD) scheme, which was developed to alleviate security assumptions on the detection devices [19, 20, 21, 22]. Then, we expand the analysis to include Continuous Variable QKD (CV-QKD) protocols [23] based on coherent states. Of particular interest are the Gaussian-modulated protocol GG02 [24] as well as a discrete-modulated approach known as QPSK [25, 26]. This is concluded with a discussion on the cost of the classical post-processing in QKD protocols. All of these protocols are detailed in Appendix C, as well as the model employed to estimate their energy consumption.
3.1 Figure of merit
@@
where the power is simply the sum of the electrical powers of the devices involved in the protocol, given in Table 3. The energy efficiency is a time-agnostic quantity which makes it an interesting figure of merit for running quantum networks. It neglects initialization costs and focuses on the energy required to achieve a certain rate with a given protocol. Note that the rate is mostly fixed by the setup and environmental parameters such as losses and noise and that inputting more energy does not necessarily increase the rate.
The second method used in this work is to fix an objective task, or metric, and study the energy required to achieve this task for different hardware, or resources, while fixing the noise (or fixing the hardware while varying the noise). In the case of QKD protocols, the objective task is fixed as the creation of $N_{\mathrm{target}}$ secret key bits between two parties. The energy required to get $N_{\mathrm{target}}$ bits of secret key with protocol and setup $\pi$ is denoted as $E^{N_{\mathrm{target}}}_{\pi}$, and can be derived from Equation 1:
@@
where $\mathcal{H}_{\pi}$ are the overall hardware elements of the protocol (including the source, the manipulation and the detection), $P_{i}$ the power of the hardware $i$ (assuming a constant consumption during the execution of the protocol), $r_{\mathrm{source}}$ is the repetition rate, and $K_{\pi}$ the secret key rate of the protocol, in bit per channel use. The execution time of the protocol is $t_{n}=\frac{N_{\rm target}}{r_{\mathrm{source}}K_{\pi}}$. Finally, $C_{\pi}^{N_{\rm target}}$ is the cost of classical computing elements, that depends on the number of target bits.
@@
$R_{\mathrm{DV}}=\left(\prod_{i}^{n}\mu_{i}\right)\left(\prod_{j}^{m}p_{j}\right)\left(\prod_{k}^{l}10^{-\eta_{k}L_{k}/10}\right),$ (4)
@@
Going from raw key to secret key in QKD protocols consists of classical rounds of communication between parties, which use part of the generated bits to assess and amplify the privacy of the rest of the bits. In this study, given the raw key rate, an estimate of the asymptotic secret key rate is given, disregarding, for now, the classical post-processing and finite size effects of these protocols. This gives a theoretical upper bound on the achievable secret key rate. Realistic implementations of QKD protocols always display some noise, quantified by the Qubit Error Rate (QBER) in DV-QKD and by the excess noise in CV-QKD. Estimating those noises is crucial to assess the quantity of secret bits that can be extracted from a string of shared bits. Appendix C explains how to derive the secret key rates $K_{\pi}$ for different protocols.
3.2 Results
@@ 3.2.1 Comparison of DV-QKD protocols
From this plot, it is clear that, for all distance regimes and using comparable hardware, BB84 is the most energy efficient QKD protocol, followed by E91 and then by MDI-QKD. With the dominating influence of detection apparatus in terms of energy consumption (see Table 3 and Appendix F), one could have expected E91 to be the worst performer due to using more detection stations. However, key rates orders of magnitude lower in MDI-QKD make it the most energy consuming, as shown in Table 2. Importantly, some qualitative advantages present in some protocols, like device independence, are not quantified in any way in this research.
@@
It should be noted that the energy efficiency, as a figure of merit, does not encompass the initialization costs of the devices. It focuses on the energy consumption of a running network, where a specific amount of power is necessary to achieve a given rate. In the inset of Figure 1, it can be seen that the result is indeed the same as what can be observed when using the protocols to generate large numbers of secret key bits. For a large $N_{\mathrm{target}}$ (here 1 Petabit), the initialization costs are absorbed in the running cost of the setup.
This result does not necessarily hold when considering a network running only for a specific task, more likely to be observed in the near-future. In Figure 2, the energy consumption $E^{\rm 1Gbit}$ of the three different DV-QKD protocols is compared for the specific task of producing 1 Gbit $=10^{9}$ bits of secret key.
@@
While the energy efficiency might be more useful in the future, when networks are constantly running, the energy required to perform a specific task gives a more precise idea of the current cost of quantum network protocols, in which initialization costs cannot be neglected.
3.2.2 Hardware study
The influence of different hardware choices on the energy cost of an implementation of the BB84 protocol can be observed in this section. The common task $E^{\rm 1Gbit}_{\rm BB84}$ was chosen to display the effects of different hardware. Equipment not considered in this study can be added to the open-source library [28].
@@ 3.2.2 Hardware study
To achieve up to 95% detection efficiency at $\lambda=1550$ nm, SNSPDs are required. They rely on cryogenic systems that are the main sources of energy consumption in a photonic setup (see Appendix F). In addition to long cooling times, which can take up to 24 hours, these cryogenic systems require large amounts of energy while they run, as they need to maintain the very low temperatures at which the detectors function. In Figure 4(a), we compare the energy consumption of an SNSPD-based setup to one using Avalanche Photodiode Detectors (APDs). APDs require much less starting time and consume less energy than SNSPDs, as they do not require cryogenics, at the cost of a detection efficiency of 25% at telecom wavelength. As shown in Figure 4(a), using APDs consumes less energy for distances up to 100 km. While the time it takes to obtain 1 Gbit of secret key is higher, the energy consumption is lower. This example illustrates a trade-off between resource cost and efficiency that can also be observed in other energy studies [4, 5]. There is an interest in choosing the energy cost as a benchmark over time: for example at 25 km, considering the same QBER, it takes on the order of 10 min to generate 1 Gbit of secret key with SNSPDs while it is around 30 min using APDs, but the energy required for the SNSPD-based setup is 60 times higher. Since the APDs can also induce more noise such as higher dark count rates, their energy consumption is also shown with a higher QBER. With a QBER of 5%, there is still a large regime where the APDs consumes less energy than the SNSPDs.
@@
(a) Comparison of the energy required to distill 1Gbit of secret key using highly efficient but energetically costly SNSPDs and less efficient, potentially noisier, but less energetically costly APDs.
@@
(b) Energy required to distill 1Gbit of secret key using different choices of wavelengths $\lambda$, in logarithmic scale.
Figure 4: Influence of hardware choices on the energetic consumption of BB84.
@@
To obtain higher detection efficiency without using cryogenic-based detectors, working at other wavelengths can be interesting. Besides the telecom range around $\lambda=1550$ nm, typical wavelength choices are near infrared ($\lambda=780$-800 nm) and visible ($\lambda=523$-532 nm). Devices for those wavelengths are included in Table 3. $E^{\rm 1Gbit}_{BB84}$ is shown in Figure 4(b) using different choices of wavelength. The critical difference between the wavelengths is the transmissivity of the fiber, as can be seen in Table 1. After 7 km of distance, working at telecom wavelengths becomes more efficient, even considering cryogenic-based detectors. It is however interesting to note that for short ($<5$ km) or very short ($<300$ m) distances, using infrared or visible wavelength can result in a lower energy consumption than the standard SNSPD-based setup at telecom wavelength. The APD-based setup at $\lambda=1550$ nm is still consuming less energy in most distance regimes, but at the cost of a lower detection efficiency. In a full-scale quantum network, it might thus be useful to consider using different wavelengths depending on the distance between all the parties involved, to optimize both the overall energy consumption of the network and the detection efficiency.
3.2.3 Continuous Variable QKD

### v2

@@ Abstract
I Introduction
@@
With the end goal of constructing a global quantum internet [1], quantum networks are rapidly developing and are already entering a deployment phase. Several national and international initiatives [2, 3] are already in place to establish an architecture allowing for distant nodes to perform quantum cryptographic tasks, such as secret key exchange. In a world with finite resources where energy demands outgrow energy generation, it is therefore crucial to estimate how much energy these networks will consume before their deployment [4]. Such a study can reveal limiting factors for future implementations of networks, or even show the energetic advantages of certain quantum technologies over classical ones. Works that estimate the energetic cost of quantum devices are few, but there are indications [5, 6] that quantum computing may show an energetic advantage before a computational one.
@@
Comparing classical and quantum communication protocols in terms of energy is a difficult task, mainly because there are no classical equivalents to most quantum networking protocols with the same level of security. For example, a quantum key distribution protocol achieving information-theoretic security cannot adequately be compared to a classical communication protocol providing only computational security. It is possible, however, to compare different quantum protocols achieving the same functionality. In this study, we focus on this task, while benchmarking the effective energy consumption of near-term quantum communication infrastructures.
@@
This work presents the foundations of a framework to estimate the energy cost of quantum network protocols. The framework is used to provide an estimation of the energy requirements of basic network functionalities, namely Quantum Key Distribution (QKD) and Conference Key Agreement (CKA), whose goals are to generate a secret private key among end users of a quantum network. These methods and hardware can be applied to most protocols based on photonic implementations, including quantum computing. In particular, the creation and sharing of entangled states among distant parties, believed to be the main goal of most quantum network architectures [7], are the building blocks of many other network protocols [8, 9, 10, 11, 12, 13].
@@
To obtain concrete figures of merit, we take a hardware-dependent approach to compare different implementations of some common protocols. Namely, we compare different QKD implementations and analyze the performance of networks of $N$ nodes, since their scaling in resources with the number of nodes is non-trivial. Using the energetic cost as a benchmark, instead of the rate or the fidelity, gives a unique perspective. For example, our simulations suggest that there exist regimes of parameters for QKD protocols where using detectors with lower detection efficiency but higher energy efficiency results in huge energy savings at the cost of increased execution time. It can also be shown that, depending on the distance, using different wavelengths can result in lower energy consumption. Moreover, we identify optimal protocols to achieve multipartite tasks as a function of the number of parties.
@@
This work is organized as follows: in Section II we introduce the model for estimating the energy consumption of photonic setups; the estimation of the energy costs of bipartite QKD protocols is presented in Section III and that of multipartite network protocols, in Section IV. To improve readability, the table of hardware components that were used as reference is deferred to Appendix A and the description of the modeling of their energy consumption, to Appendix B. Appendix C contains the description of the QKD protocols that were simulated, and additional results on time-bin based setups are presented in Appendix D, Appendix E displays experimentally measured energy values, and Appendix F touches on the distribution of the power consumption between the users.
II Energy model
@@
The energy cost of any quantum network protocol can be divided into multiple contributions, which can be studied independently. The first contribution is the source, a broad term encompassing all the required components to generate the state of light onto which the information is encoded. The second contribution comprises all manipulations required in the protocol including, for example, changes in polarization applied through motorized waveplates, phase shifters, and so on. The third is the detection, where the optical signal is measured. Finally, the last contribution comes from classical communications and computations that are required by the protocol.
@@
We define the total energy cost for a protocol $\pi$ as the time dependent function $E_{\pi}(t)$, composed of:
$\begin{split}E_{\pi}(t)&=\sum^{n_{S}}_{i}S^{i}_{\pi}(t)+\sum^{n_{M}}_{k}M^{k}_{\pi}(t)+\sum^{n_{D}}_{j}D^{j}_{\pi}(t)\\ &\hskip 14.22636pt+C_{\pi}(t)+E_{\pi}(0).\end{split}$ (1)
@@
On top of optical components, quantum network protocols typically include some classical computing elements that increase the energy cost, and those are denoted as $C_{\pi}(t)$. Two major families of contributions are included in this term. The first one consists of the classical controls of quantum components during the time of the protocol. Here, the costs of different components that are commonly involved in communication protocols are considered, such as time taggers to record timing events. All other classical costs required during a protocol are modeled by considering that an active computer is present at each node involved in the protocol. This covers classical communications and memory costs, or classical sub-protocols such as coin flipping.
The second family of contributions is the energy required for post-processing, which refers to the classical algorithms applied to the outcome of quantum processes. This cost is harder to estimate as it depends on various parameters such as the desired level of security, the classical computing architecture, and the choice of post-processing functions that are usually not standardized. In this work, this cost is mostly overlooked. However, this post-processing cost can prove to be non-trivial for some of the protocols studied in this work, especially when considering digital signal processing. This particular cost is discussed in Section III.3.
Each of these families of elements is expanded in Appendix B. In Table 3, some values are shown of power consumption for different types of devices that are used in the following sections.
III Energy cost of bipartite protocols: Quantum Key Distribution
In this section, we investigate the energetic cost of different widespread QKD protocols. Specifically, the model presented in this work is tested with the well-known BB84 protocol [14, 15]. We also explore its entanglement-based version, the E91 protocol [16, 17], along with a measurement-device-independent QKD (MDI-QKD) scheme, which was developed to alleviate security assumptions on the detection devices [18, 19, 20, 21]. Then, we expand the analysis to include Continuous Variable QKD (CV-QKD) protocols [22] based on coherent states. Of particular interest are the Gaussian-modulated protocol GG02 [23] as well as a discrete-modulated approach known as QPSK [24, 25]. This is concluded with a discussion on the cost of the classical post-processing in QKD protocols. All of these protocols are detailed in Appendix C, as well as the model employed to estimate their energy consumption.
III.1 Figure of merit
@@
where the power is simply the sum of the electrical powers of the devices involved in the protocol, given in Table 3. Energy efficiency neglects initialization costs and focuses on the energy required to achieve a certain rate with a given protocol. It is independent of the number of bits generated, which makes it an interesting figure of merit for running quantum networks. Note that the rate is mostly fixed by the setup and environmental parameters such as losses and noise and that inputting more energy does not necessarily increase the rate.
The second figure of merit used in this work is the energy required to get $N_{\mathrm{target}}$ bits of secret key with protocol and setup $\pi$. Following the Metric-Noise-Resource framework from [4], the method consists in fixing an objective task, or metric, and then study the energy required to achieve this task for different hardware, or resources, while fixing the noise (or fixing the hardware while varying the noise). In the case of QKD protocols, the objective task is fixed as the creation of $N_{\mathrm{target}}$ secret key bits between two parties. The energy cost is denoted as $E^{N_{\mathrm{target}}}_{\pi}$, and can be derived from Equation 1:
@@
where $\mathcal{H}_{\pi}$ are the overall hardware elements of the protocol (including the source, the manipulation, and the detection), $P_{i}$ the power of the hardware $i$ (assuming a constant consumption during the execution of the protocol), $r_{\mathrm{source}}$ is the repetition rate, and $K_{\pi}$ the secret key rate of the protocol, in bit per channel use. The execution time of the protocol is $t_{\pi}^{\rm target}=\frac{N_{\rm target}}{r_{\mathrm{source}}K_{\pi}}$. Finally, $C_{\pi}^{N_{\rm target}}$ is the cost of classical computing elements, that depends on the number of target bits.
@@
$R_{\mathrm{DV}}=\left(\prod_{i}^{n_{S}}\mu_{i}\right)\left(\prod_{j}p_{j}\right)\left(\prod_{k}10^{-\eta_{k}L_{k}/10}\right),$ (4)
@@
Going from raw key to secret key in QKD protocols consists of classical rounds of communication between parties, which use part of the generated bits to assess and amplify the privacy of the rest of the bits. In this study, given the raw key rate, an estimate of the asymptotic secret key rate is given, disregarding, for now, the classical post-processing and finite-size effects of these protocols. This gives a theoretical upper bound on the achievable secret key rate. Realistic implementations of QKD protocols always display some noise, quantified by the Qubit Error Rate (QBER) in DV-QKD and by the excess noise in CV-QKD. Estimating those noises is crucial to quantify the number of secret bits that can be extracted from a string of shared bits. Appendix C explains how to derive the secret key rates $K_{\pi}$ for different protocols.
III.2 Results
@@ III.2.1 Comparison of DV-QKD protocols
From this plot, it is clear that, for all distance regimes and using comparable hardware, BB84 is the most energy efficient QKD protocol, followed by E91 and then by MDI-QKD. With the dominating influence of detection apparatus in terms of energy consumption (see Table 3 and Appendix F), one could have expected E91 to be the worst performer due to using more detection stations. However, key rates orders of magnitude lower in MDI-QKD make it the most energy consuming, as shown in Table 2. Importantly, some qualitative advantages present in some protocols, like device independence, have not been considered here.
@@
It should be noted that energy efficiency, as a figure of merit, does not encompass the initialization costs of the devices. It focuses on the energy consumption of a running network, where a specific amount of power is necessary to achieve a given rate. In the inset of Figure 1, it can be seen that the result is indeed the same as what can be observed when using the protocols to generate large numbers of secret key bits. For a large $N_{\mathrm{target}}$ (here 1 Petabit), the initialization costs are absorbed in the running cost of the setup.
This result does not necessarily hold when considering a network running only for a specific task, more likely to be observed in the near future. In Figure 2, the energy consumption $E^{\rm 1Gbit}$ of the three different DV-QKD protocols is compared for the specific task of producing 1 Gbit $=10^{9}$ bits of secret key.
@@
While energy efficiency might be more useful in the future, when networks are constantly running, the energy required to perform a specific task gives a more precise idea of the current cost of quantum network protocols, in which initialization costs cannot be neglected.
III.2.2 Hardware study
The influence of different hardware choices on the energy cost of an implementation of the BB84 protocol can be observed in this section. The common task $E^{\rm 1Gbit}_{\rm BB84}$ is privileged, to display the effects of different hardware on performing the same task. Equipment not considered in this study can be added to the open-source library [27].
@@ III.2.2 Hardware study
To achieve up to 95% detection efficiency at $\lambda=1550$ nm, SNSPDs are required. They rely on cryogenic systems that are the main sources of energy consumption in a photonic setup (see Appendix F). In addition to long cooling times, which can take up to 24 hours, these cryogenic systems require large amounts of energy while they run, as they need to maintain the very low temperatures at which the detectors function. In Figure 4, we compare the energy consumption of an SNSPD-based setup to one using Avalanche Photodiode Detectors (APDs). APDs require much less starting time and consume less energy than SNSPDs, as they do not require cryogenics, at the cost of a detection efficiency of 25% at telecom wavelength. As shown in Figure 4, using APDs consumes less energy for distances up to 100 km. While the time it takes to obtain 1 Gbit of secret key is higher, the energy consumption is lower. This example illustrates a trade-off between resource cost and efficiency that can also be observed in other energy studies [4, 5]. There is an interest in choosing the energy cost as a benchmark over time: for example at 25 km, considering the same QBER, it takes on the order of 10 min to generate 1 Gbit of secret key with SNSPDs while it is around 30 min using APDs, but the energy required for the SNSPD-based setup is 60 times higher. Since the APDs can also induce more noise such as higher dark count rates, their energy consumption is also shown with a higher QBER. With a QBER of 5%, there is still a large regime where the APDs consume less energy than the SNSPDs.
The trade-off between the choice of the detector and the energy consumed to generate $N_{\rm target}$ secret key bits also depends on the length of this key. A study of the energy efficiency, that can be seen for low distances in Fig 5, shows a clear energetic advantage of using SNSPDs over APDs, for all distances. Indeed, in a running quantum network, the cost of cooling down the detector is absorbed over time and the higher rate compensates for the cost of energy. However, as we have just seen with Fig 4, there is a regime of distances and $N_{\rm target}$ where using APDs results in a lower energy cost. To illustrate this, in Fig. 4 is shown the ratio between the energy consumption of the former over the latter, for different values of the distance and $N_{\rm target}$. The green zone corresponds to a regime where using APDs results in a lower energy consumption.
[…]
@@
Figure 4: Influence of the detector choice on the energetic consumption of BB84: (a) Comparison of the energy required to distill 1Gbit of secret key using highly efficient but energetically costly SNSPDs and less efficient, potentially noisier, but less energetically costly APDs. (b) Ratio between the energy consumption of BB84 using the two different detectors, for different distances and $N_{\rm target}$. In the green zone, using APDs lead to less energy consumption.
[…]
@@
Figure 5: Influence of the wavelength choice $\lambda$ on the energetic consumption of BB84: (a) Energy required to distill 1Gbit of secret key using different choices of wavelengths, in logarithmic scale. (b)Energy efficiency of BB84 using different choices of wavelengths.
@@
To obtain higher detection efficiency without using cryogenic-based detectors, working at other wavelengths can be interesting. Besides the telecom range around $\lambda=1550$ nm, typical wavelength choices are near infrared ($\lambda=780$-800 nm) and visible ($\lambda=523$-532 nm). Devices for those wavelengths are included in Table 3. $E^{\rm 1Gbit}_{BB84}$ is shown in Figure 5 using different choices of wavelength, as well as the energy efficiency in Fig 5. The critical difference between the wavelengths is the transmissivity of the fiber, as can be seen in Table 1. After 7 km of distance, working at telecom wavelengths becomes more efficient to generate 1Gbit of secret key, even considering cryogenic-based detectors. It is however interesting to note that for short ($<5$ km) or very short ($<300$ m) distances, using infrared or visible wavelength can result in a lower energy consumption than the standard SNSPD-based setup at telecom wavelength. The APD-based setup at $\lambda=1550$ nm is still consuming less energy in most distance regimes, but at the cost of a lower detection efficiency.
The study of the energy efficiency considering different wavelengths, shown in Fig 5, shows similar results for running quantum networks. As mentioned before, at $\lambda=1550$ nm, using SNSPDs becomes more advantageous than using APDs when the network is continuously running. There is still a distance regime for which it is more advantageous to work at a different wavelength than the telecom wavelength. It is however reduced to a few hundreds of meters, which could still prove relevant in the context of the security of data-centers. In a full-scale quantum network, it might thus be useful to consider using different wavelengths depending on the distance between all the parties involved and the objective task, to optimize both the overall energy consumption of the network and the detection efficiency.
III.2.3 Continuous Variable QKD